OpenStudy (anonymous):

David kicks a soccer ball off the ground and in the air with an initial velocity of 35 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 17.9 feet 18.2 feet 18.7 feet 19.1 feet I believe it is C

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (sidsiddhartha):

H(t)= -16t2 + vt + s initially s=0 because the ball was in the ground H'(t)=-32t+35 H'(t)=0 t=35/32=1.09375 so H(1.09375)=-16(1.09375)²+35(1.09375)=19.1 so it's D hope it helps:)

OpenStudy (anonymous):

oh thank you, I got confused when it had s because i didn't know what it would be

OpenStudy (sidsiddhartha):

^_^

OpenStudy (anonymous):

V = 35 S = 0

OpenStudy (anonymous):

@sidsiddhartha Got it right, usually when it doesn't tell you what 's' is it's almost always a 0.

OpenStudy (sidsiddhartha):

yeah @greenlegodude57 :)